[figure: 9.7 cm]

HEAT CAPACITY AND ELASTIC CONSTANTS OF Fe_{3−x}Zn_{x}O_{4}

IN THE VICINITY OF THE VERWEY TRANSITION

A. Kozłowski^{1}, Z. Ka̧kol^{1}, H. Schwenk^{2}, S. Bareiter^{2}, B. Lüthi^{2}, J.M. Honig^{3}

1 Dept. of Solid State Physics, Faculty of Physics and Nuclear Techniques, University of Mining and Metallurgy, Krakw, Poland, 2 Physikalisches Institut der Universitat Frankfurt am Main, FRD, 3 Dept. of Chemistry, Purdue University, West Lafayette, IN, USA

We report on the heat capacity and elastic constants studies of Fe_{3−x}Zn_{x}O_{4} (x < 0.04) series. Adiabatic heat capacity measurements, performed on single crystals, show the clear change of the transition character from the first order for a low Zn content (x < 0.012) to higher order in the 0.012 < x < 0.035 regime. Change of the transition order is accompanied by change in the heat capacity baselines below the Verwey transition temperature T_{V}. Elastic constants measurements indicate that c_{11},(c_{11}-c_{12}) and c_{44} modes exhibit anomalies at T_{V} but only the shear wave c_{44} show critical softening while approaching the transition from high temperatures. This behavior, related to the results obtained by the heat capacity studies, point to the important role of lattice contribution to the mechanism of the Verwey transition.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : A. Kozłowski

Address for correspondence : Dept. of Solid State Physics, Faculty of Physics and Nuclear Techniques, University of Mining and Metallurgy, Krakow, Poland

Email address : kozlow@uci.agh.edu.pl